OpenStudy (anonymous):

solve -26<=3x-9<=-2 what is the solution

OpenStudy (anonymous):

Do this one like a regular equation. \[-26\le3x-9\le-2\]Add 9 to each side. \[-26+9\le3x \le-2+9\]\[-17\le3x \le7\]Divide each side by 3. \[-\frac{17}{3}\le x \le \frac{7}{3}\]

jhonyy9 (jhonyy9):

-17<=3x x>= -17/3 3x-9<=-2 3x<=7 x<=7/3